# Learning to Edit Visual Programs with Self-Supervision

### v1

@@ 4.3 Search Time
While one-shot models must author new programs from scratch without execution-feedback, our edit network has the capacity to reason over an input program, compare its execution versus the visual target, and decide how this program should be modified. As such, we hypothesize that integrating our edit network into our inference procedure will be increasingly advantageous over the OS Only approach as more time is spent on test-time search. To validate this hypothesis, we explore how the reconstruction gap between these paradigms changes as a function of time spent on search (Figure 4, left). For CSG2D we take a subset of the test-set (300 shapes) and run more rounds of our inference algorithm. As demonstrated, as more time is spent on test-time search (i.e. as the number of rounds increases) the reconstruction gap between  OS Only and OS+Edit grows wider. Moreover, we note that even on the first round there is a gap between the methods, as the one-shot network trained in the OS+Edit paradigm had access to better $P^{\text{BEST}}$entries throughout the finetuning process (i.e. the aforementioned virtuous cycle). We present qualitative results that show how the edit network evolves the population of programs towards the visual target in Figure 5.
@@ 4.4 Training with limited data
Figure 4: We compare reconstruction accuracy (Y-axis) between using an edit network and using only a one-shot network while varying time spent on inference (left) and training set size (right).
While both OS+Edit and OS Only are unsupervised in the sense that they don’t have access to any ground-truth program annotations, they do require an input set of visual data to form a target training set. We hypothesize that our edit network will be especially useful for domains with limited data (even limited unannotated data) as the program editing task is inherently more local than trying to author a complete program. Consider for instance that during finetuning, in a one-shot paradigm each visual datum can only contribute a single training example, while in our paradigm an entire distribution of edit operations can be sourced by considering the many possible edit paths one could take to transform a start program into an end program. We validate this hypothesize with an experiment where we train versions of these systems while varying the size of the target training set (Fig. 4, right). Our joint paradigm offers very strong performance even while finetuning towards an input set of just 100 training shapes (effectively matching OS Only performance with 10x more data).
@@ 4.5 Method Ablations
Edit Operations. Our default edit networks learn how to predict local edit operations from a limited set of options. We compare this paradigm with two alternatives. In the next program mode, we task the edit network with predicting all of the tokens of the program that would be created by applying the target edit operation to the input program. In the final program mode, we task the edit network with predicting the tokens of the final program associated with the visual target. This formulation was inspired by the success of denoising diffusion models for visual synthesis tasks [16] though in our setting this variant is basically an alternative one-shot model with extra conditioning information but the same target sequences. As demonstrated, neither of these approaches is as helpful as our formulation where edits are predicted as local operations. Moreover, predicting an entire program is much slower compared with predicting an edit, so fewer rounds of our inference algorithm can be run with the same time budget.
@@
Program Corruption. We source paired training data for our edit network by constructing (start, end) program pairs and then analytically finding a set of edit operations that would complete this transformation. For an alternative, we can look towards discrete diffusion methods [36, 45, 39, 29]. In our corruption variant we take inspiration from these works and design a program corruption algorithm for the Layout domain. This corruption algorithm takes an end program as input, and then samples corruption operations (i.e. inverse edit operations) that can be used as paired data for our edit network (Appendix E). As seen, this alternative formulation was not as performant as our default approach. One reason for this is that it hard to design a corruption process that converts end programs (e.g. $P^{G}$) into the distribution of programs that we have access to at inference time (e.g. $P^{S}$). Conversely, by applying our findEdits operation on $P^{G}$ and $P^{S}$ pairs, we can source paired data for our edit network that does match this distribution.
@@
Pretraining and Finetuning. In our default version there are three training phases. First, $p(z|x)$ undergoes pretraining on synthetic data. Second, $p(e|z,x)$ undergoes pretraining on synthetic data using samples from $p(z|x)$. Then both of these networks are jointly finetuned with respect to $X^{*}$. In the No FT variant, we don’t finetune either network, in no one-shot FT we don’t finetune $p(z|x)$, in no edit FT we don’t finetune $p(e|z,x)$, and in no edit PT we don’t pretrain $p(e|z,x)$. While the performance of our system remains remarkable strong even under these ablations, we get the best results by using all three training phases. Interestingly, for settings where $p(z|x)$ is not specialized for $X^{*}$, the reconstruction accuracy gap dramatically increases between the best sample in the starting population and the best sample in the final population of our inference procedure. For instance, for the no one-shot FT variant, the first round cIoU score is 0.88 which gets increased to 0.972 (0.092 improvement) through the mutations proposed by the edit model, while in our default variant the first round cIoU is 0.925 (only an improvement of .055).
@@
Inference Algorithm. We compare our inference algorithm with two alternative versions. In Naive OS we initialize the first population with $p(z|x)$, and make edits to each population member with $p(e|z,x)$, but we skip the population resampling step according to $M$ , and instead apply the highest likelihood edit from $p(e|z,x)$. While the edit network is still helpful in this paradigm (0.022 improvement from first to last round), it performs worse compared with our default implementation. In Rand+Edit, we remove $p(z|x)$ and instead fill the initial population with random program sampled from $L$. This provides a much worse initialization (0.302 cIoU in the first round), and though our edit network successfully mutates these samples towards the target, better reconstruction performance is gained by combining our edit network with an initial guesses from a one-shot model.
@@ 5 Discussion
We presented a system that learns how to edit visual programs in a goal-directed fashion. We develop a self-supervised bootstrapping approach that allows us to train an edit network for domains that lack ground-truth program annotations. We compare our proposed paradigm, that jointly finetunes a one-shot model and an edit network, against the alternative of using only a one-shot model, and find that our approach infers more accurate program reconstructions. Further, we find this performance gap is more pronounced when more time is spent on program search or when less training data is available. Finally, we justified the design of our method with an ablation experiment.
@@
While our proposed approach advances the field of visual program induction, it does come with a few limitations. Compared with prior work, we need to train another network, this impacts the time required for both pretraining and finetuning stages. Moreover, the full benefit of using an edit network is best realized with a more complex program search, and as such we use search-time budgets that are slightly more costly compared with prior work. Though our formulation would offer improved performance for work-flows that can afford to spent more time on program search, it would be useful to consider potential speed-ups of our system [7]. Finally we note that our current formulation requires access to a domain-aware findEdits operation that can analytically find a set of edits that realizes a transformation from a start program to an end program. While we find this algorithm generalizes across a range of visual programming domains, in future work, it would be interesting to consider to what degree this domain-aware procedure could be replace by more general program difference algorithms [31]. Looking ahead, we believe our framework can serve as inspiration for how to train networks that learn how to edit programs without ground-truth annotations over an even wider array of program synthesis tasks.
Appendix A Appendix
@@
Here we provide additional details on our visual programming domains (Section B), on our experimental design (Section C), on our editing operations (Section D), and on our program corruption experiments (Section E).
Appendix B Domain Details
@@
$T\xrightarrow{}Move(F,F)\ |\ Scale(F,F)\ |\ Rotate(F)\ |\ Reflect(axis)|\ P$
@@
In this domain, there are three combinator operations that combine multiple ‘shape’-typed inputs: union, difference and intersection. Move, scale, rotate and reflect are all transformation functions that consume a single ‘shape’-typed input and apply a geometric modification. Once again. Prim is a special command that produces a ‘shape’-typed argument from an input parameter variables.
@@
$T\xrightarrow{}Move(F,F,F)\ |\ Scale(F,F,F)\ |\ Rotate(F,F,F)\ |\ Reflect(axis)|\ P$
@@
As previously discussed, we follow prior work and use a synthetic pretraining phase [19, 12, 33, 37]. In this pretraining phase we randomly sample programs from the above grammars. We employ simple rejection criteria to ensure these random samples are useful (e.g. no execution errors, remain within the canvas, etc.), and find it effective to build in some of this rejection logic during the sampling phase (to improve the speed at which we can sample programs). All of the models we evaluate in our experiments train with the same sampling logic.
Appendix C Experimental Design Details
@@
For our 2D domains (2D CSG and Layout) we use a 2D-CNN. The image size of both domains is 64x64, but in 2D CSG there is only one input feature (occupancy) while in Layout there are three channels (RGB). The network we utilize consists of four layers, each containing convolution, ReLU, max-pooling, and dropout operations. Each convolution layer employs a kernel size of 3, a stride of 1, and padding of 1, with channel dimensions of 32, 64, 128, and 256 respectively. The CNN’s output is a (4x4x256) dimensional vector, which we reshape into a (16x256) vector. This vector is then processed through a 3-layer MLP with ReLU and dropout, resulting in a final (16x256) vector that serves as a 16-token encoding of the visual input. For our 3D voxel model, we adopt a similar convolutional approach by extending all 2D convolutions to 3D. We adjust the kernel size to 4, use padding of size 2, and include an additional fifth convolution layer. When processing voxel grids of size 64x64x64, this produces outputs of size (2x2x2x256). We pass these outputs through a 3-layer MLP to generate eight 256-dimensional visual tokens.
@@
We implement all of our networks in PyTorch [27]. All of our experiments are run on NVIDIA GeForce RTX 3090 graphic cards with 24GB of VRAM and consume up to 128GB of RAM (for CSG 3D experiments). We use the Adam optimizer [21] with a learning rate of 1e-4. For $p(z|x)$ pretraining we use a batch size of 128/128/64, for $p(e|z,x)$ pretraining we use a batch size of 128/128/32, for $p(z|x)$ finetuning we use a batch size of 20/20/20, and for $p(e|z,x)$ finetuning we use a batch size of 128/128/32 for Layout / 2DCSG / 3DCSG domains respectively. We pretrain on synthetic programs until convergence with respect to a validation set of synthetic program, for 34 / 17 / 18 million iterations, which takes 6 / 7 / 7 days for $p(z|x)$ and 70 / 30 / 25 million iterations, which takes 7 / 8 / 8 days for $p(e|z,x)$ for the Layout, 2D CSG, and 3D CSG domains domains. We finetune each method for a maximum of 6 days or until convergence, which took 40 / 40 / 30 bootstrap rounds for the Layout, 2D CSG and 3D CSG domains. For each finetuning run we use a $P^{G}$ set of size 10000.
@@
For our test-time inference program search we use the following population size / number of round parameters for each domain: Layout (32, 32), CSG2D (32, 32), CSG3D (80, 25). When using the Os Only method, we keep the same population / mutation general logic, but each mutation is just a randomly sampled program from $p(z|x)$. In both cases, the best reconstructing program ever seen in any round’s population is returned as the ‘chosen’ program. The settings for this method are: Layout (32, 10), CSG2D (32, 10), CSG3D (25, 25). We set these parameters so that the time spent on inference per shape is even between the two modes (5, 10, 60 seconds for the three domains). For our outer-loop inference step that populates $P^{\text{BEST}}$, we use a less expensive search time budget for both modes, approximately taking (2, 5, 10 seconds for each domain respectively). We sample programs from $p(z|x)$ with top-p (.9) nucleus sampling. We sample edits from $p(e|z,x)$ with a beam search of size 3. Interestingly, we found that this sampling strategy for Os Only outperformed a beam-search with a beam size set to the maximum number of tokens in each $L$.
Appendix D Editing Operations
As described in Section 3 our network predicts local edit operations. We find it useful to constrain the set of possible edit operations (Section 4.5).
In order to use these local edit operations, we require a few properties of the underlying DSL. We require that is a functional language, where each valid function has a ‘shape’ return type. Through a slight abuse-of-notation, we refer to functions that implicitly consume a single ‘shape’-typed argument as transformation functions (e.g. Move), and we refer to functions that consume multiple ‘shape’-typed arguments as combinator functions (e.g. Union). Note that as described in Section B, there may also be special function that instantiate ‘shape’-types from only non-‘shape’-typed arguments (e.g. Prim functions).
We allow the network to predict one of the following edit operations:

### v2

@@ 4.3 Search Time
While one-shot models must author new programs from scratch without execution-feedback, our edit network has the capacity to reason over an input program, compare its execution versus the visual target, and decide how this program should be modified. As such, we hypothesize that integrating our edit network into our inference procedure will be increasingly advantageous over the OS Only approach as more time is spent on test-time search. To validate this hypothesis, we explore how the reconstruction gap between these paradigms changes as a function of time spent on search (Figure 4, left). For 2D CSG we take a subset of the test-set (300 shapes) and run more rounds of our inference algorithm. As demonstrated, as more time is spent on test-time search (i.e. as the number of rounds increases) the reconstruction gap between  OS Only and OS+Edit grows wider. Moreover, we note that even on the first round there is a gap between the methods, as the one-shot network trained in the OS+Edit paradigm had access to better $P^{\text{BEST}}$ entries throughout the finetuning process (i.e. the aforementioned virtuous cycle). We present qualitative results that show how the edit network evolves the population of programs towards the visual target in Figure 5.
@@ 4.4 Training with limited data
Figure 4: For 2D CSG, we compare reconstruction accuracy (Chamfer distance, lower is better, Y-axis) between using an edit network and using only a one-shot network while varying time spent on inference (left) and training set size (right).
While both OS+Edit and OS Only are unsupervised in the sense that they don’t have access to any ground-truth program annotations, they do require an input set of visual data to form a target training set. We hypothesize that our edit network will be especially useful for domains with limited data (even limited unannotated data) as the program editing task is inherently more local than trying to author a complete program. Consider for instance that during finetuning, in a one-shot paradigm each visual datum can only contribute a single training example, while in our paradigm an entire distribution of edit operations can be sourced by considering the many possible edit paths one could take to transform a start program into an end program. We validate this hypothesize with an experiment where we train versions of these systems while varying the size of the target training set (Fig. 4, right). Our joint paradigm offers very strong performance even while finetuning towards an input set of just 100 training shapes, matching the performance of OS Only when it has 10x more data.
@@ 4.5 Method Ablations
Edit Operations. Our default edit networks learn how to predict local edit operations from a limited set of options. We compare this paradigm with two alternatives. In the next program mode, we task the edit network with predicting all of the tokens of the program that would be created by applying the target edit operation to the input program. In the final program mode, we task the edit network with predicting the tokens of the final program associated with the visual target. This formulation was inspired by the success of denoising diffusion models for visual synthesis tasks [16], though in our setting this variant is basically an alternative one-shot model with extra conditioning information but with the same target sequences. As demonstrated, neither of these approaches is as performant as our formulation where edits are predicted as local operations. Moreover, predicting an entire program is much slower compared with predicting an edit, so fewer rounds of our inference algorithm can be run with the same search time budget.
@@
Program Corruption. We source paired training data for our edit network by constructing (start, end) program pairs and then analytically finding a set of edit operations that would complete this transformation. For an alternative, we can look towards discrete diffusion methods [37, 46, 40, 30]. In our corruption variant we take inspiration from these works and design a program corruption algorithm for the Layout domain. This corruption algorithm takes an end program as input, and then samples corruption operations (i.e. inverse edit operations) that can be used as paired data for our edit network (Appendix F). As seen, this alternative formulation was not as performant as our default approach. One reason for this is that it hard to design a corruption process that converts end programs (e.g. $P^{G}$) into the distribution of programs that we have access to at inference time (e.g. $P^{S}$). Conversely, by applying our findEdits operation on $P^{G}$ and $P^{S}$ pairs, we can source paired data for our edit network that does match this distribution.
@@
Pretraining and Finetuning. In our default version there are three training phases. First, $p(z|x)$ undergoes pretraining on synthetic data. Second, $p(e|z,x)$ undergoes pretraining on synthetic data using samples from $p(z|x)$. Then both of these networks are jointly finetuned with respect to $X^{*}$. In the No FT variant, we don’t finetune either network, in no one-shot FT we don’t finetune $p(z|x)$, in no edit FT we don’t finetune $p(e|z,x)$, and in no edit PT we don’t pretrain $p(e|z,x)$. While the performance of our system remains remarkable strong even under these ablations, we get the best results by using all three training phases. Interestingly, for settings where $p(z|x)$ is not specialized for $X^{*}$, the reconstruction accuracy gap dramatically increases between the best sample in the starting population and the best sample in the final population of our inference procedure. For instance, for the no one-shot FT variant, the first round cIoU score is 0.88 which gets increased to 0.972 (0.092 improvement) through the mutations proposed by the edit model, while in our default variant the first round cIoU is 0.925 (an improvement of .055).
@@
Inference Algorithm. We compare our inference algorithm with two alternative versions. In Naive OS we initialize the first population with $p(z|x)$, and make edits to each population member with $p(e|z,x)$, but we skip the population resampling step according to $M$ , and instead apply the highest likelihood edit from $p(e|z,x)$. While the edit network is still helpful in this paradigm (0.022 improvement from the first to the last round), it performs worse compared with our default implementation. In Rand+Edit, we remove $p(z|x)$ and instead fill the initial population with random program sampled from $L$. This provides a much worse initialization (0.302 cIoU in the first round), and though our edit network successfully mutates these samples towards the target, better reconstruction performance is gained by combining our edit network with initial guesses from a one-shot model.
@@ 5 Discussion
We have presented a system that learns how to edit visual programs in a goal-directed fashion. We develop a self-supervised bootstrapping approach that allows us to train an edit network for domains that lack ground-truth program annotations. We compare our proposed paradigm, that jointly finetunes a one-shot model and an edit network, against the alternative of using only a one-shot model, and find that our approach infers more accurate program reconstructions. Further, we find this performance gap is more pronounced when more time is spent on program search or when less training data is available. Finally, we justified the design of our method with an ablation experiment.
@@
While our proposed approach advances the field of visual program induction, it does come with a few limitations. Compared with prior work, we need to train another network, this impacts the time required for both pretraining and finetuning stages. Moreover, the full benefit of using an edit network is best realized with a more complex program search, and as such we use search-time budgets that are slightly more costly compared with prior work. Though our formulation would offer improved performance for work-flows that can afford to spent more time on program search, it would be useful to consider potential speed-ups of our system [7]. Finally we note that our current formulation requires access to a domain-aware findEdits operation that can analytically find a set of edits that realizes a transformation from a start program to an end program. While we find that our implementation generalizes across a range of visual programming domains, in future work, it would be interesting to consider to what degree this domain-aware procedure could be replaced by more general program difference algorithms [32]. Looking ahead, we believe our framework can serve as inspiration for how to train networks that learn how to edit programs without ground-truth annotations over an even wider array of program synthesis tasks.
Acknowledgments
We would like to thank the anonymous reviewers for their helpful suggestions. Renderings of 3D shapes were produced using the Blender Cycles renderer. This work was funded in parts by NSF award #1941808 and a Brown University Presidential Fellowship. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture.
Appendix A Appendix Overview
@@
We overview the contents of our appendices. In section B we include more experimental results. We then provide additional details on our visual programming domains (Section C), on our experimental design (Section D), on our editing operations (Section E), and on our program corruption experiments (Section F).
Appendix B Additional Results
B.1 Performance on more challenging tasks
Our formulation employs a self-supervised finetuning scheme that specializes our inference networks towards a target dataset of interest. But how do our networks fare on visual inputs that are outside of these distributions? For instance, one might hypothesize that the performance gap between our joint paradigm and the one-shot paradigm might shrink when these approaches are given more challenging problems (e.g. when there is a large distribution gap between training and testing data).
Note though, that as we focus on local edits, our edit networks learn how to solve a local problem: given a current program and some visual target, we task our network with making any edit that would make the current program more similar to the target. Our hypothesis is that this framing should actually scale better than the one-shot networks when the target scenes become more complex or when they are further out-of-distribution from the training data.
Our intuition here, is that as the task complexity increases, it becomes more likely that the one-shot network will make mistakes. The edit network is able to account for the mistakes of the one-shot network and suggest local fixes that make improvements in a goal-directed fashion. When the target is out-of-distribution, even if the edit network has not seen a similar example, it can still compare the current program’s execution against the target scene. Reasoning over the differences between the two states admits a more local task (as evidenced by our data efficient learning), and this property can aid in generalization.
To validate the above hypothesis, we set up an experiment to compare how our formulation (which uses a one-shot and edit network jointly) performs against using only the one-shot network for more challenging tasks in the Layout and 3D CSG domains. For the Layout domain, we evaluate the methods on scenes of new “challenge” concepts (e.g. butterflies / snowmen) that were not seen in the training / validation sets. For 3D CSG, we evaluate the methods on “challenge” shapes from other categories of ShapeNet (airplanes, knives, lamps, laptops, motorbikes, mugs, pistols, skateboards, rifles, vessels) that were not part of the original finetuning training set (chairs, tables, benches, couches).
Table 3: We evaluate reconstruction accuracy for "challenge" tasks that come from concepts or categories not present in the target training set. For both layout and 3D CSG, we observe that our joint paradigm that integrates an edit network with one-shot models outperforms the alternative of using only one-shot models.
Layout cIoU $\Uparrow$ 3D CSG IoU $\Uparrow$
OS Only 75.8 60.8
OS + Edit 87.6 70.9
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
OS Only OS + Edit (Ours) Target
Figure 6: Qualitative reconstructions of "challenge" tasks for 3D CSG.
Using the same models from Section 4.2, we compare the reconstruction performance for these challenge tasks. In Table 3, we report the reconstruction performance over 192 challenge tasks for the layout domain and 100 challenge tasks for the 3D CSG domain. As seen from both the quantitative and qualitative comparisons (Figures  6 and  7), it’s clear that our approach, which utilizes both the one-shot and edit networks, outperforms using only the one-shot network for these more challenging program induction tasks, even when they are further outside the training distribution.
GPT 4V GPT 4V (ICE) OS + GPT 4V OS Only OS + Edit (Ours) Target
Figure 7: Qualitative reconstructions of "challenge" tasks for the layout domain. We compare against GPT-4V in a zero-shot setting (column 1), when an in-content example (ICE) is provided in the prompt (column 2), and when the one-shot model’s predicted program is provided as input (column 3). Our approach (column 5) finds more accurate reconstructions of these out-of-distribution targets (column 6) compared with using only the one-shot network (column 4).
B.2 Comparison to large vision-language models
As discussed in Section 2, there has been much recent research that has investigated how LLMs can aid in program synthesis tasks. Relatedly, some works have even begun to examine to what extent large vision-language models are able to understand programs that capture visual data [28]. Following similar ideas, we ran an experiment to explore how well large vision-language models (e.g. GPT-4v) are able to perform on our visual program induction tasks.
We provide some qualitative results of using GPT-4v to predict visual programs on examples from our layout domain in Figure 7. These predictions were made with a relatively straightforward prompt containing: a task-description, a description of the DSL, and the input image that should be reconstructed (zero-shot, col 1). We then tried improving this prompt by adding an in-context example of a (program, image) pair (one-shot, col 2). We also experimented with providing GPT-4v with a program predicted from the one-shot network, along with this program’s execution, and asking it to edit the program to make it more similar to the target image (col 3). Please find full prompts for these experiments at https://github.com/rkjones4/VPI-Edit .
As can be seen, GPT-4v in this setting proved inferior to our proposed method (col 5). While we do not include these results to say that these sorts of large vision-language models will not ever be of use for this task, we do believe that these results showcase that this task is not easily solved with currently available frontier models.
B.3 Method Ablations on 2D CSG domain
Table 4: Ablation study on our method for the 2D CSG domain.
Method Chamfer Distance $\Downarrow$
Ours (default) 0.111
No FT 0.321
No one-shot FT 0.230
No edit FT 0.123
No edit PT 0.145
In Section 4.5 we presented results for an ablation experiment on the layout domain. We include additional ablation results on the 2D CSG domain in Table 4. Note that while some ablation conditions do come close to our default performance (e.g. no edit FT) these ablation conditions are also made possible by our contributions, as they all use an edit network. When comparing our method against an alternative without an edit network (OS Only, Table 1) we have consistently seen that our method offers a meaningful improvement. Below we offer some additional commentary on these results.
No edit FT
In this ablation condition the edit network is pretrained (with synthetic random data), but is then kept frozen during the joint finetuning. As the task of the edit network is mostly local, we find that the edit network is able to achieve impressive performance even when it does not get to finetune towards data in the target distribution. That said, the edit network is still very important in this ablation condition (if it’s removed then this condition becomes OS Only). Even though the edit network remains fixed during finetuning, it still helps to find better solutions during inner-loop inference (Alg 1, line 5), and this better training data leads to a better one-shot network. However, once again, the performance of the system is maximized when the edit network is also allowed to update during finetuning.
No one-shot FT
This condition does impressively well for the layout domain. This is because even though the one-shot network is much worse in this setting, the edit network can overcome almost all of its mistakes, as layout is a relatively easier domain. Consider that for the layout domain, the default approach has a starting cIoU of 0.925 (initialized from the one-shot network, which is finetuned) which gets improved to 0.980 through improvements made by the edit network. However, the one-shot network of this ablation condition drops the starting cIoU to 0.88 (when it is kept frozen), and yet the edit network is still able to raise this performance all the way to 0.972 (explaining the strong reconstruction score of this condition). That said, when considering the 2D CSG ablation results in Table 4, we see that for more complex domains it is critical to also finetune the one-shot network, as this ablation condition achieves only a Chamfer distance of 0.230 compared with the Chamfer distance of 0.111 achieved by our default approach.
Appendix C Domain Details
@@
$T\xrightarrow{}Move(F,F)\ |\ Scale(F,F)\ |\ Rotate(F)\ |\ Reflect(axis)$
@@
In this domain, there are three combinator operations that combine multiple ‘shape’-typed inputs: union, difference and intersection. Move, scale, rotate and reflect are all transformation functions that consume a single ‘shape’-typed input and apply a geometric modification. Once again, Prim is a special command that produces a ‘shape’-typed argument from only a parameter-type argument.
@@
$T\xrightarrow{}Move(F,F,F)\ |\ Scale(F,F,F)\ |\ Rotate(F,F,F)\ |\ Reflect(axis)$
@@
As previously discussed, we follow prior work and use a synthetic pretraining phase [19, 12, 34, 38]. In this pretraining phase we randomly sample programs from the above grammars. We employ simple rejection criteria to ensure these random samples are useful (e.g. no execution errors, outputs remain within the canvas, etc.), and find it effective to build in some of this rejection logic during the sampling phase (to improve the speed at which we can sample programs). All of the models we evaluate in our experiments train with the same sampling logic.
Appendix D Experimental Design Details
@@
For our 2D domains (2D CSG and Layout) we use a 2D CNN. The image size of both domains is 64x64, but in 2D CSG there is only one input feature (occupancy) while in Layout there are three channels (RGB). The network we utilize consists of four layers, each containing convolution, ReLU, max-pooling, and dropout operations. Each convolution layer employs a kernel size of 3, a stride of 1, and padding of 1, with channel dimensions of 32, 64, 128, and 256 respectively. The CNN’s output is a (4x4x256) dimensional vector, which we reshape into a (16x256) vector. This vector is then processed through a 3-layer MLP with ReLU and dropout, resulting in a final (16x256) vector that serves as a 16-token encoding of the visual input. For our 3D CNN model, we adopt a similar convolutional approach by extending all 2D convolutions to 3D. We adjust the kernel size to 4, use padding of size 2. When processing voxel grids of size $32^{3}$, this produces outputs of size (2x2x2x256). We pass these outputs through a 3-layer MLP to generate eight 256-dimensional visual tokens.
@@
We implement all of our networks in PyTorch [27]. All of our experiments are run on NVIDIA GeForce RTX 3090 graphic cards with 24GB of VRAM and consume up to 128GB of RAM (for 3D CSG experiments). We use the Adam optimizer [21] with a learning rate of 1e-4. For $p(z|x)$ pretraining we use a batch size of 128/128/64, for $p(e|z,x)$ pretraining we use a batch size of 128/128/32, for $p(z|x)$ finetuning we use a batch size of 20/20/20, and for $p(e|z,x)$ finetuning we use a batch size of 128/128/32 for Layout / 2D CSG / 3D CSG domains respectively. We pretrain on synthetic programs until convergence with respect to a validation set of synthetic program, for 34 / 17 / 18 million iterations, which takes 6 / 7 / 7 days for $p(z|x)$ and 70 / 30 / 25 million iterations, which takes 7 / 8 / 8 days for $p(e|z,x)$ for the Layout, 2D CSG, and 3D CSG domains respectively. We finetune each method for a maximum of 6 days or until convergence, which took 40 / 40 / 30 bootstrap rounds for the Layout, 2D CSG and 3D CSG domains. For each finetuning run we use a $P^{G}$ set of size 10000.
@@
For our test-time inference program search we use the following population size / number of round parameters for each domain: Layout (32, 32), 2D CSG (32, 32), 3D CSG (80, 25). When using the Os Only method, we keep the same population / mutation general logic, but each mutation is just a randomly sampled program from $p(z|x)$. In both cases, the best reconstructing program ever seen in any round’s population is returned as the ‘chosen’ program. The settings for this method are: Layout (32, 10), 2D CSG (32, 10), 3D CSG (25, 25). We set these parameters so that the time spent on inference per shape is even between the two modes (5, 10, 60 seconds for the three domains). For our inner-loop inference step that populates $P^{\text{BEST}}$, we use a less expensive search time budget for both modes, approximately taking (2, 5, 10 seconds for each domain respectively). We sample programs from $p(z|x)$ with top-p (.9) nucleus sampling. We sample edits from $p(e|z,x)$ with a beam search of size 3. Interestingly, we found that this sampling strategy for Os Only outperformed a beam-search with a beam size set to the maximum number of tokens in each $L$.
Appendix E Visual Program Edits
E.1 Local Edit Operations
As described in Section 3, our network predicts local edit operations. We find it useful to constrain the set of possible edit operations as described in Section 4.5.
In order to use these local edit operations, we require a few properties of the underlying DSL. We require that it is a functional language, where each valid function has a ‘shape’ return type. Through a slight abuse-of-notation, we refer to functions that implicitly consume a single ‘shape’-typed argument as transformation functions (e.g. Move), and we refer to functions that consume multiple ‘shape’-typed arguments as combinator functions (e.g. Union). Note that as described in Section C, there may also be special functions that instantiate ‘shape’-types from only non-‘shape’-typed arguments (e.g. Prim functions).
Specifically, our formulation allows the network to predict one of the following edit operations: